OpenStudy (anonymous):

if you ignore friction, the amount of time required for a box to slide down an inclined plane is the square root of 2a/g sin theta cos theta, where a is the horizontal distance defined by the inclined plane, g is the acceleration due to gravity, and theta is the angle of the inclined plane. for what values of theta is the expression undefined?

OpenStudy (zzr0ck3r):

\[friction=\frac{2a}{g*\sin(\theta)}\]?

OpenStudy (anonymous):

2a over g times sin theta times cos theta

OpenStudy (zzr0ck3r):

\[friction=\frac{2a}{g*\sin(\theta)cos(\theta)}\]?

OpenStudy (anonymous):

Yes

OpenStudy (zzr0ck3r):

so its undefined when sin(theta) = 0 or cos(theta) = 0 so when \[\theta=\frac{\pi}{2}\pm\pi*k\\and\\\theta=0\pm\pi*k\]

OpenStudy (zzr0ck3r):

so when \[\theta=0\pm\frac{\pi}{2}\]

OpenStudy (anonymous):

So it would equal 0 degrees?

OpenStudy (zzr0ck3r):

um there are infinite theta for which it is undefined

OpenStudy (anonymous):

Oh

OpenStudy (zzr0ck3r):

\[\{....,-2pi,-1.5\pi,-\pi,-.5\pi,0,0.5\pi,pi,1.5\pi.2\pi......\}\]

OpenStudy (zzr0ck3r):

\[\theta=\pm\frac{\pi}{2}\]

OpenStudy (anonymous):

Thank you, I think I got it now(: